OpenStudy (anonymous):

If the % yield is 82.0% how many grams of silicon is needed to make 105g of chrominium by the reaction? 3SI(s) + 2 Cr2O3 (S)--> 3 SiO2 (s) + 4 Cr(I)

OpenStudy (anonymous):

% yield = actual / theoretical You were given the actual. You can find the ratio of the theoretical. Make sense? :-)

OpenStudy (anonymous):

4 moles of Chromium are in the products, 3 moles of Silicon in the reactants

OpenStudy (anonymous):

\[(0.82) = \frac{105g }{theoretical}\]

OpenStudy (anonymous):

ok, so all I have to do is use the equation backward, since I am given the percentyield

OpenStudy (anonymous):

Yeah basically, you'll have the theoretical/optimal mass of chromium, then you need to use the stoichiometric ratios from the balanced equation reaction.

OpenStudy (anonymous):

They're giving you the inefficiency %, you just need to find out how much you originally had to put in to get at least 105g of that particular product.

OpenStudy (anonymous):

This is actually pretty real-world industry stuff ;-)

OpenStudy (anonymous):

oh, ok that seem to make sense.

OpenStudy (anonymous):

Your boss will typically say this is how much we need to produce per hour or per day or whatever and you'll know roughly how inefficient your equipment is. And then you find the equipment is getting extra inefficient one day so you go do a maintenance check. :-P

OpenStudy (anonymous):

or finding the % yield given testing, you know what to expect for output

OpenStudy (anonymous):

It's not just a chemistry concept I what I'm trying to get at.

OpenStudy (anonymous):

could you assist me on solving it. Want to make sure I got the right answer and solved it propperly

OpenStudy (anonymous):

Go ahead, post your work and I'll be happy to check it for accuracy ;-) See that "\(\Sigma\) Equation" button below? It's handy :-D type in frac{numerator}{denominator} while in it to get something like this: \[\frac{numerator}{denominator} \rightarrow stuff\]

OpenStudy (anonymous):

rightarrow turns into \(\rightarrow\)

OpenStudy (anonymous):

The other two conversion factors you'll need are the "amu" (aka. molecular weights) for Silicon and Chromium which are in grams per mole.

OpenStudy (anonymous):

Ok. to find the theoretical Yield of Cr, do i need to solve it like thi.s Theoretical yield of Cr = 100/82 x 105 = 128.05g (5sf)

OpenStudy (anonymous):

Professional tip, don't convert to decimals unless you have to, because it introduces error: 105g Cr / 82% =\(\frac{5250}{41} g Cr \approx\) 128g Cr \[\large \frac{5250}{41} g_{Cr} \cdot \frac{1 mol_{Cr}}{51.996 g_{Cr}} \cdot \frac{3 mol_{Si}}{4 mol_{Cr}} \cdot \frac{28.085 g_{Si}}{1 mol_{Si}} = ?\] Always write your units, it'll help make sure you're doing things right. The units all cancel out here to give you an answer in...?

OpenStudy (anonymous):

Your final answer will have 3 significant figures btw, if you're wondering about rounding.

OpenStudy (anonymous):

Make sense? :-)

OpenStudy (anonymous):

yea, thanks